OpenStudy (anonymous):

If x^2 is divisible by 4 then x is eve. How are we supposed to prove this implication in Discrete Mathematics? Using truth tables? or by some other way?

OpenStudy (anonymous):

*even

ganeshie8 (ganeshie8):

try proving the contrapositive

OpenStudy (anonymous):

how?

ganeshie8 (ganeshie8):

whats the contrapositive of given conditional statement ?

OpenStudy (anonymous):

If x is not even the x^2 is not divisible by 4

ganeshie8 (ganeshie8):

yes, `not even` can be replaced by `odd`

ganeshie8 (ganeshie8):

statement : `if x^2 is divisible by 4, then x is even` contrapositive : `if x id odd, then x^2 is not divisible by 4`

ganeshie8 (ganeshie8):

since the original statement and contrapositive follow the same truth values, proving contrapositive is same as proving the original statement

ganeshie8 (ganeshie8):

`if x id odd, then x^2 is not divisible by 4` Proof : say x = 2k + 1 x^2 = (2k+1)^2 = 4k^2 + 4k + 1 = 2q+1 = odd number an odd number is not divisible by 4

OpenStudy (anonymous):

If we are given the postulate that p implies q = not q implies not p and q implies p = not p implies not q

OpenStudy (anonymous):

then can we use truth tables to prove these two and conclude that since these are true our initial implication is true. would that be correct?

ganeshie8 (ganeshie8):

is that another question ?

OpenStudy (anonymous):

no no. I'm asking if we do it like this would that be correct. I actually asked can this be proved using truth tables in any way?

ganeshie8 (ganeshie8):

never seen anything like that before, but my knowledge in math so tiny so...

OpenStudy (anonymous):

like this :p

ganeshie8 (ganeshie8):

nice :) but why are you doing both : 1) p->q and 2) q->p ?

ganeshie8 (ganeshie8):

you just need to prove it in one direction right ?

OpenStudy (anonymous):

Yea. just for the sake of it. Because we did both in class and the assignment is related to that.

OpenStudy (amistre64):

how does showing that an original statements truth value is equal to its contrapositive; and the converse of the original is equal to its own contrapositive .... prove that the original statement is true? the original statement if the cow is alive, then the cow is sleeping has the same truth value as: if the cow is not sleeping, then the cow is not alive. the converse statement: if the cow is sleeping, then the cow is alive. has the same truth value as : if the cow is not alive then the cow is not sleeping i cant see how this proves the original statement, without knowing if P and Q share the same truth value to start with. is there an error in my thinking?

OpenStudy (anonymous):

I'm not sure I know the reason. I just know that that is how we prove an impilcation and that if the contra-positive is true, the implication is true. It's a rule or something.

OpenStudy (amistre64):

but how do we show that its true to start with?

OpenStudy (amistre64):

in a lot of statements, its easier to show that the contrapositive is true, rather than the original statements original structure.

OpenStudy (anonymous):

I honestly have no idea. I'm just a student. I wouldn't be asking if I knew :p

OpenStudy (amistre64):

If x^2 is divisible by 4 then x is even contraP If x is odd, then x^2 is not divisible by 4 let x be an odd number: x = 2n+1 square it: (2n+1)^2 = 4n^2 +2n + 1 factor out a 4: 4 (n^2 +n/2 + 1/4) and when we divide by 4, we get: n^2 +n/2 + 1/4 showing that the results is never an integer .... just a thought

OpenStudy (anonymous):

how does that prove our implication or it's contra-positive?

OpenStudy (amistre64):

trying to recall if there is a suitable way to approach it from there ....

OpenStudy (anonymous):

There might be. But it eludes me :p

OpenStudy (amistre64):

4n^2 +2n + 1 = 4k, for some integer k 4n^2 +2n + 1-4k = 0 is a quadratic, that we can solve for n \[n=\frac{-2\pm\sqrt{4-4(4)(1-4k)}}{2(4)}\] \[n=\frac{-2\pm\sqrt{4(1-4(1-4k))}}{2(4)}\] \[n=\frac{-2\pm2\sqrt{1-4+16k}}{2(4)}\] \[n=\frac{-1\pm\sqrt{16k-3}}{4}\] got to her elol

OpenStudy (amistre64):

16k-3 >= 0 16k >= 3 k>= 3/16 so k has to be at best be an integer that is greater or equal to 1 but im prolly overthinking stuff ...

OpenStudy (anonymous):

yea you are.

OpenStudy (amistre64):

:) good luck with it then

OpenStudy (anonymous):

Thanks. Thankyou for your time too :)

OpenStudy (amistre64):

If x^2 is divisible by 4 then x is even contraP If x is odd, then x^2 is not divisible by 4 let x be an odd number: x = 2n+1, for some integer n square it: (2n+1)^2 = 4n^2 +2n + 1 this is where i started overthinking it .... so to get it back on track: factor it like this: 2(2n^2 +n) + 1 now n^2 + n is an integer since the set of integers is closed under addition and multiplication let k = n^2 + n to clean this up to see that the results of squaring is just another odd number: 2k + 1 if it is divisible by 4 then for some integer, m ... 2k + 1 = 4m 2k - 4m + 1 = 0 2(k-2m) + 1 = 0 again because of closure, k-2m is just some integer and what we have shown is that in order for the square of an odd number to be divisible by 4, then 0 must be an odd number. 0 is not an odd number, so we have proven our contrapositive to be true, namely: If x is odd, then x^2 is not divisible by 4 is true